# Conditional Panoramic Image Generation via Masked Autoregressive Modeling

### v1

@@ Abstract
Recent progress in panoramic image generation has underscored two critical limitations in existing approaches. First, most methods are built upon diffusion models, which are inherently ill-suited for equirectangular projection (ERP) panoramas due to the violation of the identically and independently distributed (i.i.d.) Gaussian noise assumption caused by their spherical mapping. Second, these methods often treat text-conditioned generation (text-to-panorama) and image-conditioned generation (panorama outpainting) as separate tasks, relying on distinct architectures and task-specific data. In this work, we propose a unified framework, Panoramic AutoRegressive model (PAR), which leverages masked autoregressive modeling to address these challenges. PAR avoids the i.i.d. assumption constraint and integrates text and image conditioning into a cohesive architecture, enabling seamless generation across tasks. To address the inherent discontinuity in existing generative models, we introduce circular padding to enhance spatial coherence and propose a consistency alignment strategy to improve generation quality. Extensive experiments demonstrate competitive performance in text-to-image generation and panorama outpainting tasks while showcasing promising scalability and generalization capabilities. Project Page: https://wang-chaoyang.github.io/project/par.
@@ 1 Introduction
Despite some progress in recent years, existing panorama generation methods [17, 54, 55, 71, 67, 26, 35, 59, 75, 62, 61, 74] predominantly rely on diffusion-based base models and face two significant problems: task separation and inherent limitations of the modeling approach. First, different panoramic generation tasks are highly bound to specific foundation models. For instance, most text-to-panorama models [17, 54, 55, 71, 67] are based on Stable Diffusion (SD) [46, 41] while panorama outpainting models [26, 35, 59] are fine-tuned on SD-inpainting variants, resulting in task-specific pipelines that lack flexibility. Second, diffusion-based models struggle to handle the unique characteristics of equirectangular projection (ERP) panoramas, which map spherical data onto a 2D plane, violating the i.i.d. Gaussian noise assumption.
In addition to these fundamental issues, redundant modeling in existing methods introduces unnecessary computational costs. This redundancy manifests in several ways: 1) Some approaches [28, 35] iteratively inpaint and warp local areas, leading to error accumulation over successive steps. 2) Others [71, 53] adopt a dual-branch structure, with global and local branches, where the local branches are not directly used as final output targets. These challenges highlight the lack of a simple, unified, and efficient paradigm for conditional panoramic image generation.
In this work, we explore a powerful yet much less explored panoramic image generation paradigm to address the above limitations - the violation of i.i.d. assumption and the lack of unification. Moreover, we allow the model to take text or images as input and control the generated content at any position on the canvas. As shown in Fig. 1, PAR is competent for text-to-panorama, panorama outpainting, and editing tasks. Notably, our PAR requires no task-specific data engineering - both conditions are seamlessly integrated via a single likelihood objective. In comparison, Omni2 [65] relies on meticulously designed joint training datasets to align heterogeneous tasks, inevitably introducing complexity and domain bias.
@@
Our contribution can be summarized as follows: 1) We propose PAR, which fundamentally avoids the conflict between ERP and i.i.d. assumption raised by diffusion models, and also seamlessly integrates text and image-conditioned generation within a single architecture. 2) To enforce better adaptation to the panorama characteristic, we propose specialized designs, including dual-space circular padding and a translation consistency loss. 3) We evaluate our model on popular benchmarks, demonstrating competitive performance in text and image-conditional panoramic image generation tasks and showcasing its scalability and generalization capabilities.
@@ 2 Related Work
Conditional Panoramic Image Generation. There are two primary task paradigms in this area: 1) text-to-panorama (T2P), which synthesizes 360-degree scenes from textual descriptions, and 2) panorama outpainting (PO), which extends partial visual inputs to full panoramas. Given the success of diffusion models (DMs) [30, 31, 32, 41, 46, 49, 50] on generation tasks, such as image generation [46, 72, 22, 41, 40, 16], image editing [3, 36, 47, 73], image super resolution [21, 48, 68], video generation [2, 19, 66], current approaches typically build upon DMs, yet face two problems. First, DM is inherently unsuitable for panorama tasks due to reliance on the i.i.d. Gaussian noise assumption, which ERP transformations violate. Second, existing approaches lack task unification. Specifically, T2P methods typically fine-tune SD, while PO solutions adapt the SD-inpainting variants. Although some work [65] attempts a unified solution, it is based on DM and relies on complex task-specific data engineering. From a methodological perspective, existing methods can be divided into bottom-up and top-down. The former generates images with limited FoV sequentially [28, 35] or simultaneously [55, 26] and finally stitches them together, but faces the problem of lacking global information and error accumulation. The latter [71, 53] usually adopts a dual-branch architecture including a panorama and a perspective branch. However, this design brings unnecessary computational cost. In parallel, there are several AR-based attempts. They either rely on large DMs for final generation [35, 18], are limited to low resolution with additional up-sampling modules [9], or stitch perspective images without correct panorama geometry [77]. Overall, existing methods do not provide a unified and theoretically sound solution for conditional panorama generation, which we rethink through MAR modeling.
@@
Autoregressive Modeling. Motivated by the success of large language models (LLMs) [4, 56], AR models, with great scalability and generalizability, are becoming a strong challenger to diffusion models. Several studies delve into visual AR with two distinct pipelines. One pipeline adopts raster-scan modeling [51, 45, 14, 15, 69, 60, 64, 27, 38], which treats the image as visual sentences and sequentially predicts the next token. However, this strategy does not fully exploit the spatial relationships of images. To make up for the shortcomings of raster-scan modeling, some works [6, 23, 70, 57, 29, 13] extend the concept of masked generative models [7] to AR modeling, namely MAR. Instead of predicting one token within one forward, the model randomly selects several candidate positions and predicts attending to the decoded positions. Intuitively, this modeling provides flexible control over arbitrary spatial positions simply by adjusting the order of tokens to be decoded. Hence, we resort to MAR as a flexible way to unify text and image-conditioned panorama generation while avoiding the theoretical defects of panorama DMs.
@@ 3.1 Preliminaries
Conditional Autoregressive Models formulate conditional data generation as a sequential process through next token prediction. Given an ordered sequence of tokens $\{{x}_{t}\}_{t=1}^{T}$, the joint distribution $p({x|c})$ factorizes into a product of conditional probabilities:
@@ 3.3 Masked Autoregressive Modeling with Consistency Alignment
Unlike the standard AR modeling, the decoder’s output is not used as the final result, but as a conditional signal $z$ to drive a lightweight denoising network, which is implemented by a three-layer multilayer perceptron (MLP) $\epsilon_{\theta}$. During the training phase, the noise-corrupted $x_{t}$ is denoised by $\epsilon_{\theta}$ under $z$ as follows:
@@ 4 Experiment
Modeling Method ##\##params T2P PO PE FAED $\downarrow$ FID $\downarrow$ CS $\uparrow$ DS $\downarrow$
@@ 4.1 Experiment Settings
Several metrics are used in this paper. Fréchet Inception Distance (FID) [20] measures the similarity between the distribution of real and generated images in a feature space derived from a pre-trained inception network. Lower FID scores indicate better realism of images. CLIP Score (CS) [44] evaluates the alignment between text and image. Furthermore, as FID relies on an Inception network [52] trained on perspective images, we also report the Fréchet Auto-Encoder Distance (FAED) [39] score, which is a variant of FID customized for panorama. To measure the cycle consistency of panoramic images, we adopt Discontinuity Score (DS) [10].
@@
Table 2: Panorama outpainting results on the Matterport3D dataset. For FID-h, we horizontally sample 8 perspective images with FoV=$90\degree$ for each panorama and then calculate their FID.
@@ 4.2 Main Results
Text-to-Panorama. Tab. 1 shows the quantitative comparison results. Within the scope of AR-based methods, our PAR significantly outperforms previous approaches regarding generation quality and continuity score. Our method is slightly inferior in the CS metric, probably because CLIP [44] is pre-trained on massive perspective images, and panoramic geometry is not well aligned with perspective images. Moreover, our model performs decently compared with a strong diffusion-based baseline, PanFusion, with only 0.3B parameters. Specifically, PAR-0.3B outperforms PanFusion with 1.73 and 4.06 on FAED and FID, with a discontinuity score of 0.58. Scaling up the parameters helps achieve better results, which is further illustrated in Sec. 4.3.
@@ 4.3 Ablation Study
Generalization. Fig. 6 visualizes the outpainting results of PAR on OOD data. SUN360 [63] is used for evaluation, which has different data distributions from Matterport3D. PAR shows decent performance across various scenarios. We also compared with several PO baselines trained on the SUN360 dataset. However, Diffusion360 [17] suffers from realistic scenes and lacks details. Panodiff [59] also encounters artifact problems (red sky in the $1^{st}$ row).
@@
Method FID $\downarrow$ CS $\uparrow$ DS $\downarrow$
@@ 5 Conclusion
This paper proposes PAR, a MAR-based architecture unifying text and image-conditioned panoramic image generation. We delve into the cyclic translation equivariance of ERP panoramas and correspondingly propose a consistency alignment and circular padding strategy. Experiments demonstrate PAR’s effectiveness on text-to-panorama and panorama outpainting tasks, whilst showing promising scalability and generalization capability. We hope our research provides a new solution for the panorama generation community.

### v2

@@ Abstract
Recent progress in panoramic image generation has underscored two critical limitations in existing approaches. First, most methods are built upon diffusion models, which are inherently ill-suited for equirectangular projection (ERP) panoramas due to the violation of the identically and independently distributed (i.i.d.) Gaussian noise assumption caused by their spherical mapping. Second, these methods often treat text-conditioned generation (text-to-panorama) and image-conditioned generation (panorama outpainting) as separate tasks, relying on distinct architectures and task-specific data. In this work, we propose a unified framework, Panoramic AutoRegressive model (PAR), which leverages masked autoregressive modeling to address these challenges. PAR avoids the i.i.d. assumption constraint and integrates text and image conditioning into a cohesive architecture, enabling seamless generation across tasks. To address the inherent discontinuity in existing generative models, we introduce circular padding to enhance spatial coherence and propose a consistency alignment strategy to improve the generation quality. Extensive experiments demonstrate competitive performance in text-to-image generation and panorama outpainting tasks while showcasing promising scalability and generalization capabilities. Project Page: https://wang-chaoyang.github.io/project/par.
@@ 1 Introduction
Despite some progress in recent years, existing panorama generation methods [18, 56, 57, 74, 70, 27, 36, 62, 77, 65, 64, 76] predominantly rely on diffusion-based base models and face two significant problems: task separation and inherent limitations of the modeling approach. First, different panoramic generation tasks are highly bound to specific foundation models. For instance, most text-to-panorama models [18, 56, 57, 74, 70] are based on Stable Diffusion (SD) [48, 43]. In contrast, panorama outpainting models [27, 36, 62] are fine-tuned on SD-inpainting variants, resulting in task-specific pipelines that lack flexibility. Second, diffusion-based models struggle to handle the unique characteristics of equirectangular projection (ERP) panoramas, which map spherical data onto a 2D plane, violating the i.i.d. Gaussian noise assumption.
In addition to these fundamental issues, existing methods also contain redundant modeling, which introduces unnecessary computational costs. This redundancy manifests in several ways: 1) Some approaches [29, 36] iteratively inpaint and warp local areas, leading to error accumulation over successive steps. 2) Others [74, 55] adopt a dual-branch structure, with global and local branches, where the local branches are not directly used as final output targets. These challenges underscore the need for a straightforward, unified, and efficient paradigm for generating conditional panoramic images.
In this work, we explore a powerful yet less explored panoramic image generation paradigm to address the above limitations: the violation of the i.i.d. assumption and the lack of unification. Moreover, we allow the model to take text or images as input and control the generated content at any position on the canvas. As shown in Fig. 1, PAR is competent for text-to-panorama, panorama outpainting, and editing tasks. Notably, our PAR requires no task-specific data engineering - both conditions are seamlessly integrated via a single likelihood objective. In comparison, Omni2 [68] relies on meticulously designed joint training datasets to align heterogeneous tasks, inevitably introducing complexity and domain bias.
@@
Our contribution can be summarized as follows: 1) We propose PAR, which fundamentally avoids the conflict between ERP and i.i.d. assumption raised by diffusion models, and also seamlessly integrates text and image-conditioned generation within a single architecture. 2) To enforce better adaptation to the panorama characteristic, we propose specialized designs, including dual-space circular padding and a translation consistency loss. 3) We evaluate our model on popular benchmarks, demonstrating competitive performance in text and image-conditional panoramic image generation tasks. The results also prove the scalability and generalization capabilities of our model.
@@ 2 Related Work
Conditional Panoramic Image Generation. There are two primary task paradigms in this area: 1) text-to-panorama (T2P), which synthesizes 360-degree scenes from textual descriptions, and 2) panorama outpainting (PO), which extends partial visual inputs to full panoramas. Given the success of diffusion models (DMs) [31, 32, 33, 43, 48, 51, 52] on generation tasks, such as image generation [48, 75, 23, 43, 42, 17], image editing [4, 37, 49], image super resolution [22, 50, 71], video generation [3, 20, 69], current approaches typically build upon DMs, yet face two problems. First, DM is inherently unsuitable for panorama tasks due to reliance on the i.i.d. Gaussian noise assumption, which ERP transformations violate. Second, existing approaches lack task unification. Specifically, T2P methods typically fine-tune SD, while PO solutions adapt the SD-inpainting variants. Although some work [68] attempts a unified solution, it is based on DM and relies on complex task-specific data engineering. From a methodological perspective, existing methods can be categorized into two main approaches: bottom-up and top-down. The former generates images with limited FoV sequentially [29, 36] or simultaneously [57, 27] and finally stitches them together, but faces the problem of lacking global information and error accumulation. The latter [74, 55] typically employs a dual-branch architecture comprising a panorama branch and a perspective branch. However, this design brings unnecessary computational cost. In parallel, several AR-based attempts have been made. They either rely on large DMs for final generation [36, 19], are limited to low resolution with additional up-sampling modules [10], or stitch perspective images without correct panorama geometry [80]. Overall, existing methods do not provide a unified and theoretically sound solution for conditional panorama generation, which we rethink through MAR modeling.
@@
Autoregressive Modeling. Motivated by the success of large language models (LLMs) [5, 58], AR models, with great scalability and generalizability, are becoming a strong challenger to diffusion models. Several studies delve into visual AR with two distinct pipelines. One pipeline adopts raster-scan modeling [53, 47, 15, 16, 72, 63, 67, 28, 39], which treats the image as visual sentences and sequentially predicts the next token. However, this strategy does not fully exploit the spatial relationships of images. To make up for the shortcomings of raster-scan modeling, some works [7, 24, 73, 59, 30, 14, 1] extend the concept of masked generative models [8] to AR modeling, namely MAR. Instead of predicting one token within one forward, the model randomly selects several candidate positions and predicts attending to the decoded positions. Intuitively, this modeling provides flexible control over arbitrary spatial positions simply by adjusting the order of tokens to be decoded. Hence, we resort to MAR as a flexible way to unify text and image-conditioned panorama generation while avoiding the theoretical defects of panorama DMs.
@@ 3.1 Preliminaries
Conditional Autoregressive Models formulate conditional data generation as a sequential process through next token prediction. Given an ordered sequence of tokens $\{{x}_{t}\}_{t=1}^{T}$, the joint distribution p(x|c)p({x|c}) factorizes into a product of conditional probabilities:
@@ 3.3 Masked Autoregressive Modeling with Consistency Alignment
Unlike the standard AR modeling, the decoder’s output is not used as the final result, but as a conditional signal $z$ to drive a lightweight denoising network, which is implemented by a multilayer perceptron (MLP) $\epsilon_{\theta}$. During the training phase, the noise-corrupted $x_{t}$ is denoised by $\epsilon_{\theta}$ under $z$ as follows:
@@ 4 Experiment
Modeling Method #\#params T2P PO PE FAED $\downarrow$ FID $\downarrow$ CLIP Score $\uparrow$ DS $\downarrow$
@@ 4.1 Experiment Settings
Several metrics are used in this paper. Fréchet Inception Distance (FID) [21] measures the similarity between the distribution of real and generated images in a feature space derived from a pre-trained inception network. Lower FID scores indicate better realism of synthesized results. CLIP Score (CS) [46] evaluates the alignment between text and image. Furthermore, as FID relies on an Inception network [54] trained on perspective images, we also report the Fréchet Auto-Encoder Distance (FAED) [41] score, which is a variant of FID customized for panorama. To measure the cycle consistency of panoramic images, we adopt Discontinuity Score (DS) [11].
@@
[…]
@@ 4.2 Main Results
Text-to-Panorama. Tab. 1 shows the quantitative comparison results. Within the scope of AR-based methods, our PAR significantly outperforms previous approaches regarding generation quality and continuity score. Our method is slightly inferior in the CLIP score metric, probably because CLIP [46] is pre-trained on massive perspective images, and panoramic geometry is not well aligned with perspective images. Moreover, our model performs decently compared with a strong diffusion-based baseline, PanFusion, with only 0.3B parameters. Specifically, PAR-0.3B outperforms PanFusion with 1.73 and 4.06 on FAED and FID, with a discontinuity score of 0.58. Scaling up the parameters helps achieve better results, which is further illustrated in Sec. 4.3.
@@ 4.3 Ablation Study
Generalization. Fig. 6 visualizes the outpainting results of PAR on OOD data. SUN360 [66] is used for evaluation, which has different data distributions from Matterport3D. PAR shows decent performance across various scenarios. We also compared with several PO baselines trained on the SUN360 dataset. However, Diffusion360 [18] suffers from unrealistic scenes and lacks details. Panodiff [62] also encounters artifact problems (red sky in the $1^{st}$ row).
@@
Method FID $\downarrow$ CLIP Score $\uparrow$ DS $\downarrow$
@@ 5 Conclusion
This paper proposes PAR, a MAR-based architecture that unifies text- and image-conditioned panoramic image generation. We delve into the cyclic translation equivariance of ERP panoramas and correspondingly propose a consistency alignment and circular padding strategy. Experiments demonstrate PAR’s effectiveness on text-to-panorama and panorama outpainting tasks, whilst showing promising scalability and generalization capability. We hope our research provides a new solution for the panorama generation community.
Acknowledgement. This work was supported by the National Key Research and Development Program of China (No. 2023YFC3807600).
Appendix A Proof of Non-i.i.d. of ERP Transformation
A.1 Problem Setting
Let the unit sphere
$\mathbb{S}^{2}=\{(x,y,z)\in\mathbb{R}^{3}\,:\,x^{2}+y^{2}+z^{2}=1\}$
be parameterised by spherical angles $(\phi,\theta)\in(-\pi,\pi]\times[0,\pi].$ The equirectangular projection of spatial resolution $(H,W)\in\mathbb{N}^{2}$ is the mapping
$T:[0,W)\times[0,H)\longrightarrow(-\pi,\pi]\times[0,\pi],\qquad(u,v)\mapsto\Bigl(\phi(u),\theta(v)\Bigr)$
with
$\phi(u)=\frac{2\pi u}{W}-\pi,\qquad\theta(v)=\frac{\pi v}{H}.$
Conversely, $u=\dfrac{\phi+\pi}{2\pi}W,\;v=\dfrac{\theta}{\pi}H.$
A.2 Gaussian noise field on the sphere
Let $\{\,\varepsilon(\phi,\theta)\,\}_{(\phi,\theta)\in(-\pi,\pi]\times[0,\pi]}$ be a Gaussian noise field on the sphere, i.e.
$\mathbb{E}[\varepsilon(\phi,\theta)]=0,\qquad\mathrm{Cov}\!\bigl(\varepsilon(\phi,\theta),\varepsilon(\phi^{\prime},\theta^{\prime})\bigr)=\sigma^{2}\,\delta(\phi-\phi^{\prime})\,\delta(\theta-\theta^{\prime}),$ (10)
where $\delta$ is the Dirac distribution and $\sigma^{2}>0$ is a constant variance density (per steradian).
A.3 Pixel value definition in ERP
Each ERP pixel $(u,v)\in\{0,\dots,W-1\}\times\{0,\dots,H-1\}$ covers a spherical patch
$P_{uv}\;=\;\Bigl[\phi(u-\tfrac{1}{2}),\phi(u+\tfrac{1}{2})\Bigr)\times\Bigl[\theta(v-\tfrac{1}{2}),\theta(v+\tfrac{1}{2})\Bigr).$
Its area is
$A_{uv}=\int_{P_{uv}}\!\sin\theta\,d\phi\,d\theta\;=\;\frac{2\pi}{W}\;\bigl[\cos\theta(v-\tfrac{1}{2})-\cos\theta(v+\tfrac{1}{2})\bigr]\;=\;\frac{2\pi^{2}}{WH}\;\sin\theta_{v},$ (11)
with the mid‐latitude $\theta_{v}:=\theta(v)=\pi v/H$.
The noise value stored in the pixel is the area–average
$\eta_{uv}:=\frac{1}{A_{uv}}\int_{P_{uv}}\varepsilon(\phi,\theta)\,dA,\quad dA=\sin\theta\,d\phi\,d\theta.$ (12)
Because the integral in Eq. 12 is a linear functional of the Gaussian field $\varepsilon$, each $\eta_{uv}$ is still Gaussian.
A.4 Independence
For two different pixels $(u,v)\neq(u^{\prime},v^{\prime})$ the patches $P_{uv},P_{u^{\prime}v^{\prime}}$ are disjoint. Since Eq. 10 states that $\varepsilon(\phi,\theta)$ is white, integrals over disjoint regions are uncorrelated, hence Gaussian independent:
$\mathrm{Cov}(\eta_{uv},\eta_{u^{\prime}v^{\prime}})=\iint_{P_{uv}}\!\iint_{P_{u^{\prime}v^{\prime}}}\!\frac{\sigma^{2}\,\delta(\phi-\phi^{\prime})\delta(\theta-\theta^{\prime})}{A_{uv}A_{u^{\prime}v^{\prime}}}\,dA\,dA^{\prime}=0\quad\Longrightarrow\quad\eta_{uv}\;\perp\!\!\!\perp\;\eta_{u^{\prime}v^{\prime}}.$ (13)
A.5 Non-identical distribution
The variance of one pixel, using Eq. 10, Eq. 11, and Eq. 12, is
$\mathrm{Var}(\eta_{uv})=\frac{1}{A_{uv}^{2}}\iint_{P_{uv}}\!\iint_{P_{uv}}\sigma^{2}\,\delta(\phi-\phi^{\prime})\delta(\theta-\theta^{\prime})\,dA\,dA^{\prime}=\frac{\sigma^{2}}{A_{uv}}=\sigma^{2}\;\frac{WH}{2\pi^{2}}\;\frac{1}{\sin\theta_{v}}.$ (14)
Eq. 14 depends on the latitude index $v$ through the factor $\sin\theta_{v}$. Pixels close to the poles ($\theta_{v}\approx 0$ or $\pi$) possess a larger variance than pixels near the equator ($\theta_{v}\approx\pi/2$). Hence
$\mathrm{Var}(\eta_{uv})\neq\mathrm{Var}(\eta_{u,v^{\prime}})\quad\text{for }v\neq v^{\prime}.$
Appendix B More Details on Method
Implementation Details. The sampling step for PAR is set to 64 and $r=0.125$ by default. We use Janus Pro-7B to get the captions with the template as “Use one sentence to describe the photo in detail.". We then calculate the token lengths and keep them less than 77. Experiments shows that less than $0.1\%$ of the captions do not meet the requirements, and we add "and briefly" at the end of those prompts.
Framework Details. The transformer takes patchified visual tokens (from the VAE encoder), masking indicators, and text embeddings as inputs. It outputs a conditional signal to drive the subsequent denoising network MLP. The decoding mechanism is illustrated in Fig. 9.
The MLP has multiple stacked blocks. Within each block, the input is first processed by AdaLNZero to get gating coefficients. It is then transformed using a projection layer consisting of two linear layers with silu activation. The result is then LN-normalized and multiplied by the gating coefficients, and finally combined with a residual connection.
Figure 9: Illustration of decoding mechanism.
Appendix C More Empirical Studies and Visual Results
Time Analysis. It takes about 2 days to train the 1.4B model on 8 NVIDIA A100 GPUs. As shown in Tab. 4, we benchmark the inference speed for the 0.3B model with different AR steps using one NVIDIA A100 GPU with a batch size of 8. The denoising steps for MLP are 25.
Table 4: Ablations of inference time under different AR steps.
AR steps 16 32 64
Inference speed (sec/image) $\downarrow$ 3.02 5.49 10.03
We also compare PAR-0.3B with PanFusion with 64 AR steps in Tab. 5.
Table 5: Comparison with PanFusion regarding inference time.
PanFusion PAR (ours)
Inference speed (sec/image) $\downarrow$ 28.91 10.03
FID $\downarrow$ 45.21 41.15
As shown in Tab. 6, circular padding does not bring about additional computational cost in the transformer and MLP, which account for the majority of inference time. The AR step is 16.
Table 6: Ablations of different padding ratios regarding inference time.
Padding Ratio 0 0.0625 0.125 0.25 0.5
Inference speed (sec/image) $\downarrow$ 2.99 3.01 3.02 3.04 3.05
Additional costs 0 $0.67\%$ $1.00\%$ $1.67\%$ $2.01\%$
Table 7: Text-to-panorama results on the Matterport3D dataset.
FAED $\downarrow$ FID $\downarrow$ CLIP Score $\uparrow$ DS $\downarrow$
DiffPano [70] 10.03 53.29 30.31 6.16
UniPano [40] 5.87 44.74 30.45 0.77
PAR-0.3B (ours) 3.39 41.15 30.21 0.58
More Results on Text-to-Panorama. Tab. 7 shows that our method has advantages on text-to-panorama task. We also provide more visual results in Fig. 10. Our model can generate reasonable panoramic images under the guidance of textual prompts. Tab. 9 provides ablation study for different denoising steps with our 0.3B model.
Figure 10: Visualization of text-to-panorama generation, including the generated results and corresponding textual prompts.
Table 8: Panorama outpainting results on the Matterport3D dataset.
FID $\downarrow$ FID-h $\downarrow$
PanoDiff [62] 65.11 46.58
PAR w/o prompt 41.63 25.97
Table 9: Ablations of denoising steps.
Steps 10 20 30 40
FID $\downarrow$ 41.57 40.19 41.98 43.77
More Results on Panorama Outpainting. Tab. 8 shows that our method has advantages on panorama outpainting task. We provide more visual results for panorama outpainting in Fig. 11. We test the performance of our model in indoor setting and outdoor setting, and our model can achieve decent results, which proves its generalizability.
Figure 11: Visualization of panorama outpainting results.
Visual Results of Circular Padding. Fig. 12 compares different circular padding strategies, including $r_{pre}=0$ and $r_{pre}=0.5$. $r_{post}$ ranges from 0 to 0.5. It converges quickly with $r_{post}>0$. $r_{pre}$ helps to maintain cycle consistency at the semantic level.
Figure 12: Visualization results of circular padding under different settings.
Classifier-free Guidance. We evaluate PAR-1.4B with different CFG coefficients. The model obtains 40.04 FID when CFG=3 and 39.76 FID when CFG = 10, indicating that too large or too small guidance strength deteriorates the quality of the generation.
Figure 13: Perspective projections of synthesized results on Matterport3D dataset.
Ablations on other datasets. We include additional experiments using the Structured3D [78] dataset, which is a large-scale synthesized indoor dataset for house design with well-preserved poles. We sampled 9000 images for training and 1000 for testing. In Tab. 10, our method outperforms PanoLlama. Fig. 14 visualizes the synthesized results of PAR-0.3B. This indicates that the blur in the pole regions originates from the dataset and is irrelevant to the model, which can also be demonstrated by the visualizations of perspective projections in Fig. 13.
Table 10: Comparisons of PAR-0.3B and PanoLlama on Structured3D dataset.
FID $\downarrow$ CLIP Score $\uparrow$ DS $\downarrow$
PanoLlama [80] 125.35 33.90 16.37
PAR-0.3B 47.02 30.75 0.68
Figure 14: Visualizations of synthesized results on Structured3D dataset.
Zero-shot inference on OOD datasets. We evaluate panorama discontinuity using both our method and StitchDiffusion [61]. StitchDiffusion forces the model to fuse seam areas during inference. On the contrary, our pre-padding makes the model learn semantic consistency during the training phase, and post-padding achieves pixel-level consistency in the inference stage. In this experiment, PAR-0.3B and StitchDiffusion generate images for 100 text prompts sampled from a subset of SUN360, respectively. Our method achieves a DS score of 0.63, outperforming StitchDiffusion’s 1.12.
For zero-shot outpainting, we compare the two methods on 100 images from the Structured3D dataset, which is excluded from both training sets for fairness. As shown in Tab. 11, our model outperforms Diffusion360.
Table 11: Zero-shot panorama outpainting results.
FID $\downarrow$ FID-h $\downarrow$
Diffusion360 [18] 140.91 74.89
PAR w/o prompt 127.01 68.27
Failure case. As shown in Fig. 15, our model may fail in some details of small objects, such as tables and sofas.
Figure 15: Failure cases.
Appendix D Discussion
Comparisons to multi-view diffusion models. On the one hand, our AR-based method has several advantages over multi-view diffusion models (MVDMs). First, Computational Efficiency. MVDMs require generating several perspective images to cover a panorama, resulting in more computational cost due to the attention mechanism. Second, Global Consistency. Coordinating multiple views demands complex attention mechanisms to ensure coherence, and struggles to maintain a consistent global understanding - sometimes leading to duplicated or inconsistent content. Third, Fewer Seam Issues. MVDMs must handle more seams at the boundaries between views, which increases the difficulty of achieving seamless stitching and can introduce visible artifacts.
On the other hand, since our model directly generates the entire panoramic image, it needs to use panoramic data to drive the pre-trained model (trained on perspective images) to adapt to the panoramic domain knowledge, which might be less efficient.
Broader Impacts. Our proposed PAR employs autoregressive set-by-set modeling for panoramic image generation tasks. It has significant advantages in generation performance compared with previous methods. Moreover, it shows the capability of zero-shot task generalization, like panorama outpainting and editing. This is an attempt towards the unification model in the field of panoramic image generation and will motivate people to explore the unification of different tasks in panorama generation. From the perspective of social impact, this will help the creation of artistic content, but at the same time it may result in the problem of fake content.
Safeguards. The content generated by the generative model, including AR models in this paper, demonstrates a high degree of correlation with the training data. From this perspective, it is important to ensure the fairness and cleanliness of the training data. In this way, we can effectively prevent the model from generating harmful content.
Limitations. We find that there is still a gap between the generated results and the real panoramic images, such as some details and textures. We argue that scaling on more realistic images can help alleviate this problem. However, due to the scarcity of panoramic data, we leave it as future work.